the bulk modulus of a metal is 8*10^9 n/m^2 and its density is 11 gm per cubic cm the density of this metal under a pressure of 20 000 n/cm^2 willbe

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
From the given datas,
Bulk modulus = 8*10^9 n/m^2
Density = 11 gm/cm^3
Pressure = 20000 n/cm^2
Density = Bulk modulus * Density /pressure
              = 8 * 10^9 * 11 /2 * 10^ 4
              = 44 * 10^5 g/cm^3.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)